how do you solve for these equations 4x+2y=9 & 2x+10y=9 using elimination technique because I don't understand.
Answer by TimothyLamb(4379)

4x + 2y = 9
2x + 10y = 9
---
multiply by -2:
2x + 10y = 9
-4x - 20y = -18
---
add:
4x + 2y = 9
-4x - 20y = -18
---
-18y = -9
---
y = 1/2
x = ?
---
4x + 2y = 9
4x = 9 - 2y
x = 9/4 - (1/2)y
---
x = 9/4 - (1/2)(1/2)
x = 9/4 - 1/4
x = 8/4
x = 2
---
Answer:
x = 2
y = 1/2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php